Use the compound interest formulas A=p(1+r/n)^nt and A=pe^rt to solve.
Find the accumulated value of an investment of $1130 at 8% compounded annually for 14 years.

Answer by Boreal(15235)

The first is A=1130(1+.08)^14; 1.08^14=2.937. Without rounding, multiply by $1130=$3319.03
The second is 1130*e^(0.08*14); e^1.12=3.065, and result is $3463.29